?? ?? m_{gauche}·d_{gauche} = m_{droite}·d_{droite}

?? m_{gauche} m_{droite}

\frac{0, 1}{1} = sin(α_{support}) = \frac{0, 2}{2}

A A_{gauche} = 2·9, 81·0, 1 = 1·9, 81·0, 2 = A_{droite}

A = F · d A A_{F,d} = F ·d

A F d ??

A_{F,d} = \overrightarrow{F ·}\overrightarrow{d} = ∥\overrightarrow{F}∥·∥\overrightarrow{d}∥ cos(α) = F ·d·cos(α)

?? \overrightarrow{F} \overrightarrow{d}

\overrightarrow{F} ↑↑ \overrightarrow{d}

\overrightarrow{F} \overrightarrow{d}

\overrightarrow{F}⊥\overrightarrow{d} cos(α) =

0 \overrightarrow{F} \overrightarrow{d}

\overrightarrow{F} ↑↓ \overrightarrow{d}

??

\overrightarrow{Δl} A_{i}

\overrightarrow{F_{i}} \overrightarrow{Δl_{i}}

A_{i} = \overrightarrow{F_{i}}·\overrightarrow{Δl_{i}}

A_{i}

A_{AB} ≅ ∑_{i=1}^{n} A_{i} = ∑_{i=1}^{n} \overrightarrow{F}·\overrightarrow{Δl_{i}}

\overrightarrow{Δl_{i}} \overrightarrow{Δl_{i}}

A_{AB} = \begin{matrix}lim \\ \overrightarrow{Δl_{i}} → 0 \end{matrix} ∑_{i}^{∞} \overrightarrow{F}·\overrightarrow{Δl_{i}} = ∫_{A}^{B} \overrightarrow{F}·\overrightarrow{dl}

A = ∫_{A}^{B} \overrightarrow{F}·\overrightarrow{dl}

[A] = [F]·[l] = N·m = J = Joules

A = F ·d = 5·5 = 25

A = F ·d·cos(α) = 5·5·cos(20) = 23, 5

A = F ·d·cos(α) = 5·5·cos(180) = −25

l

\begin{matrix}A & = ∫_{0}^{5} \overrightarrow{F ·}\overrightarrow{dl}\begin{matrix}= ∫_{0}^{5} F · dl \\ \begin{matrix}↑ \\ car \end{matrix} & \overrightarrow{F} & ↑↑ & \overrightarrow{dl} \end{matrix} = ∫_{0}^{5} l · dl \\ = \frac{1}{2} · l^{2}|_{0}^{5} = \frac{1}{2} · (25 − 0) = 12, 5 J \end{matrix}

m h m h

\overrightarrow{F}·\overrightarrow{d} = m\overrightarrow{g}·

\overrightarrow{h} = mg·

hcar m\overrightarrow{g} ↑↑ \overrightarrow{h} = mg·

(h_{i}− h_{f}) = mgh_{i}−

mgh_{f} = E_{pot i}−

E_{pot f} = −ΔE_{pot}

h h_{i}− h_{f}

E_{pot} = m·g·h

m v_{o} v

\overrightarrow{F}·\overrightarrow{d} = F ·

d\overrightarrow{F} ↑↑ \overrightarrow{d} = ma· dF = m· a = m· \frac{v^{2}−v_{o}^{2}}{2·d} ·